# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 1. $\cdot$
    Computing excluded minors. A major challenge at the interface of computer science and graph-minor theory is to determine the excluded minors for explicit minor-closed graph-classes. Partial solutions to 1.1 for $k\leqslant 3$ can be used to determine the excluded minors for planar graphs and series-parallel graphs. We have conviction that solutions for larger $k$ imply exact structure theorems for other graph-classes, such as the class of apex-graphs, and that these structure theorems can be used to determine the excluded minors of these classes.
@@ 2. $\cdot$
    Connectivity augmentation. Carmesin and Sridharan used the canonicity and explicitness of the tri-separation decomposition [Tridecomp] to construct the first FPT-algorithm for connectivity augmentation from 0 to 4 [conny_aug]. We would be excited to see if Theorem 1 can be used to extend this to 5-connectivity.
@@ 4. $\cdot$
    Parallel computing. Total-nestedness makes computing the decomposition in parallel a feasible option, as all partial solutions can be merged without conflict; see for example [Sarah].
  5. $\cdot$
    Cops and robber. In the cops and robber game of Robertson and Seymour, $k$ cops have a winning strategy for catching a robber in a graph iff that graph has a tree-decomposition of width $<k$ [GMcopsRobber]. When the robber is invisible, however, the probability of the cops catching the robber depends on the minimum number of leaf-bags attained by any tree-decomposition of width $<k$. While with other approaches to constructing small-width tree-decompositions, it is not even clear how to verify that they minimize the number of leaf-bags, our theorem gives such tree-decompositions right away in the case of $k$ cops searching $k$-connected graphs for $k\leqslant 4$ [CopsInvisibleRobber].
  6. $\cdot$
    Groups. The characterisation of vertex-transitivity for 4-connectivity (2) crucially relies on canonicity and total-nestedness. In a different direction, local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [StallingsNilpotent]. The proof heavily relies on the Tutte-type decomposition for local 2-separators [Local2sep] and exploits the combination of canonicity and total-nestedness. A next step could be using Theorem 1 and 2, or the tri-separation decomposition [Tridecomp].
Comparison to the tri-separation decomposition. Carmesin and Kurkofka introduced the first Tutte-type canonical decomposition for 3-connected graphs [Tridecomp]. They used tri-separations instead of strict tri-separations to obtain their decomposition. The main difference between tri-separations on the one hand, and strict tri-separations and tetra-separations on the other hand, is how their degree-conditions handle neighbours in the separator: tri-separations count them, but strict tri-separations and tetra-separations ignore them. This seemingly subtle difference has a surprisingly big impact: the verbatim extension of the tri-separation approach to 4-connectivity fails 2.
For 3-connected graphs, the strict tri-separation decomposition can be viewed as a refinement of the tri-separation decomposition, but the details are a bit technical; see Appendix B.
@@ 1.1. Overview of the proof
To get started, let $G$ be a 4-connected graph. Let $N$ denote its set of totally-nested tetra-separations, and let $\mathcal{T}=(T,\mathcal{V})$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by $N$. Let $\tau$ denote the torso of $\mathcal{T}$ at an arbitrary node $t$ of $T$. It will be straightforward to check that $\tau$ is 4-connected or a $K_{4}$.
In order to show that $\tau$ satisfies one of the outcomes of Theorem 1, we will study how $\tau$ interacts with the tetra-separations of $G$. Indeed, the outcome will be determined by this. Let $(A,B)$ be a tetra-separation of $G$, which may possibly be crossed. We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$. Here may think of $(A,B)$ as ‘cutting through’ $\tau$. Whether $\tau$ is interlaced or not already yields a first outcome:
@@
(See 10.5.) A technical lemma from [Tridecomp] does most of the work of proving 1.
By 1, we may assume from now on that some tetra-separation $(A,B)$ of $G$ interlaces $\tau$. Then $(A,B)$ is not in $N$, so some tetra-separation $(C,D)$ crosses $(A,B)$. The remaining outcomes will be determined by the way in which $(A,B)$ and $(C,D)$ cross. More specifically, the outcome will be determined by the size of the vertex-centre $Z$, which consists of all vertices in the intersection of the separators of $(A,B)$ and $(C,D)$. We will show that
@@
(See the (Crossing Lemma). (4.1).) So there will be exactly three outcomes. Here we focus only on the cycle-of-graphs outcome, which we get for $|Z|=0$. This outcome is the hardest to prove, and we use all the new methods introduced in this paper in the proof of this outcome. It is also different from all situations encountered in [Tridecomp].
Assume now that $|Z|=0$. As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure 1.12. Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions. We will then merge the four Tutte-decompositions into a cycle-decomposition $\mathcal{O}=(O,(G_{o})_{o\in V(O)})$. So all adhesion-sets of $\mathcal{O}$ have size two. We then use the 4-connectivity of $G$ to note that the torsos of $\mathcal{O}$ are triangles, 4-cycles, or 3-connected. Moreover, we will find some rules on how these types of torsos may be arranged on $O$. For instance, a torso that is a 4-cycle must always have two 3-connected torsos as neighbours, and if two triangle-torsos are neighbours then their adhesion-set is spanned by an edge.
@@
(See 9.41). In order to show the hard backward implication of 3, we will introduce a
@@
(See 6.16.) Examples of connectivity-properties of a tetra-separator $S(A,B)$ that are necessary for $(A,B)$ to be totally-nested include, roughly speaking, that for every bipartition $(\pi_{1},\pi_{2})$ of $S(A,B)$ into two classes $\pi_{i}$ of size 2 there must be
@@
A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure 1.13 for an example.
@@
With 4 at hand, the problem of proving the hard backward implication of 3 becomes a problem about finding many paths in many configurations, roughly speaking. To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove 3. Finding the paths takes a lot of work, especially when $O$ is short; see Figure 1.14 for example.
Figure 1.14. The bag $G_{o}$ is the top left quarter. Its candidate $(U_{o},W_{o})$ for a totally-nested tetra-separation is indicated in yellow. The sets $\pi_{1}$ and $\pi_{2}$ consist of the blue and the green vertices, respectively. As a small part of proving 3 via 4, we have to find 5 independent $\pi_{1}$$\pi_{2}$ paths (red). We find two such paths in $G_{o}$ using Menger as $G_{o}$ is good. To find the other three, we essentially go around $O$ in both directions starting from $o$, constructing one path in the one direction and two in the other. The example shows that the construction must be highly efficient to avoid collision of the paths when the decomposition-cycle $O$ is short.
By 3, each tetra-separation $(U_{o},W_{o})$ for a good torso $H_{o}$ corresponds to a directed edge $e_{o}$ of $T$. The second key result about $\mathcal{O}$ will be that
@@
where $t$ is the node of $T$ with $\tau$ as torso. Phrased differently, the tetra-separations $(U_{o},W_{o})$ from good torsos $H_{o}$ are precisely the elements of $N$ that ‘carve out’ $\tau$. (See Lemma 9.49.)
Once we have proved 5, we are almost done. Indeed, it now follows that the torso $\tau$ can be obtained from $G$ by essentially replacing the good torsos of $\mathcal{O}$ with $K_{4}$’s and contracting a few edges. Then it will be fairly immediate to see that $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. (See 9.50.)
Proving 5, however, is a little tricky. We will assume for a contradicting that 5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$. Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure 1.15. We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$. Eventually, there will be essentially only one bag $G_{o}$ left, and we will see that its torso $H_{o}$ is good. We will then show that $(E,F)=(U_{o},W_{o})$ is the only possibility. Since $(U_{o},W_{o})$ does not ‘cut through’ $\mathcal{O}$, this will yield a contradiction.
@@
Outlook. We would be excited to see an extension of Theorem 1 to higher connectivity. However, the verbatim extension of tetra-separations fails 2 for $k=5$.
Open Problem 1.4.
Extend Theorem 1 to $k$-connectivity for $k\geqslant 5$.
Both the strict tri-separation decomposition and the tri-separation decomposition may have quasi-4-connected torsos. To decompose these torsos further, we can apply the tetra-separation decomposition, but only to those quasi-4-connected torsos that are even 4-connected.
Open Problem 1.5.
Extend Theorem 1 from 4-connectivity to quasi-4-connectivity.
Simply using the tetra-separation decomposition fails: see Figure 1.16 for an illustrative example that essentially generalises 2 to quasi-4-connectivity.
Figure 1.16. A quasi-4-connected version of 2: all tetra-separations are crossed, yet the graph is far from $5$-connected and each $K_{10}$ can be replaced with an arbitrary 5-connected graph
The tri-separation decomposition has been used to prove a characterisation of vertex-transitivity for 3-connectivity [Tridecomp, Corollary 2]. One of the outcomes is that the graph is quasi-4-connected or a triangle-expansion of a quasi-4-connected graph. [^2] The actual outcome is slightly more specific.
Open Problem 1.6.
Extend 2 from 4-connectivity to quasi-4-connectivity.
We believe that a solution to 1.5 can be used to approach 1.6.
Open Problem 1.7.
Extend the FPT-algorithm of [conny_aug] for connectivity augmentation from 0 to 4 to 5-connectivity.
Since the tri-separation decomposition [Tridecomp] has been heavily exploited in [conny_aug], we hope that Theorem 1 can be used to approach 1.7.
@@
For directed graphs, the case $k=1$ of 1.1 has recently been solved almost completely by Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [Flo].
@@
Organisation of the paper. Section 2 introduces or recalls all essential terminology. Section 3 provides stronger but more technical versions of Theorem 1 and 2, and proves all while assuming a number of key lemmata that remain to be proved later in the paper. Section 4 provides a characterisation of how tetra-separations cross. Section 5 introduces a systematic way for obtaining tetra-separations from mixed-4-separations. Section 6 gives a characterisation of total-nestedness for tetra-separations in terms of connectivity-properties of their separators. The next sections investigate the different outcomes of Theorem 1: Section 7 deals with the $K_{4,m}$; Section 8 deals with the double-wheel; Section 9 deals with the cycle-of-graphs; and Section 10 deals with quasi-5-connectivity. This will conclude the proof of Theorem 1. Section 11 introduces an ‘angry theorem’ for 4-connectivity: a characterisation of the 4-connected graphs all whose tetra-separations are crossed (hence the name). Section 12 uses Theorem 1 to obtain a characterisation of the vertex-transitive 4-connected finite graphs. Appendix A proves that the circular saws from 1 really exhibit the properties we used. Appendix B compares the strict tri-separation-decomposition (2) with the tri-separation decomposition [Tridecomp].
@@ 2. Essential terminology
This section introduces all terminology used by Theorem 1. For graph-theoretic terminology we follow [bibel].
@@ 2.1. Tetra-separations
A mixed-separation+ of a graph $G$ is a pair $(A,B)$ such that $A\cup B=V(G)$. The separator $)𝑆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{S(A,B)}}$ of $(A,B)$ is the disjoint union of the vertex set $A\cap B$ and the edge-set $E(A\smallsetminus B,B\smallsetminus A)$. The order of $(A,B)$ is the size of $S(A,B)$. A mixed-separation+ of order $k$ is also called mixed-$k$-separation+. A mixed-separation+ is proper if both $A\smallsetminus B$ and $B\smallsetminus A$ are nonempty. In this case, we refer to $(A,B)$ as a mixed-separation without the plus-sign. A mixed-separation of order $k$ is also called mixed-$k$-separation. A mixed-separation with $k$ vertices in its separator but no edges is called a separation, or $k$-separation.
@@
  1. degree-condition:
    every vertex in $A\cap B$ has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
  2. matching-condition:
    the edges in $S(A,B)$ form a matching.
@@
Recall that a graph $G$ is quasi-$k$-connected if it is $(k-1)$-connected and every $(k-1)$-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
@@
We will prove Lemma 2.3 in Section 5.2.
@@
The lower-bound $|G|\geqslant 8$ in Lemma 2.3 (ii) is tight: the graph depicted in Figure 2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.

### v2

@@ 1. $\cdot$
    Connectivity augmentation. Carmesin and Sridharan used the canonicity and explicitness of the tri-separation decomposition [CK23] to construct the first FPT-algorithm for connectivity augmentation from 0 to 4 [CS25\$ˆ+\$]. That is, they showed that there is an algorithm with runtime $C(\ell)\cdot n^{\mathcal{O}(1)}$ and the following specifications:
    Input  : a graph $G$ on $n$ vertices, a number $\ell\in\mathbb{N}$ and a set $F\subseteq[V(G)]^{2}\smallsetminus E(G)$ of non-edges. Output  : a set $X\subseteq F$ of $\leqslant\ell$ edges such that $G+X$ is 4-connected, if such an $X$ exists, and ‘no’ otherwise. We strongly believe that Theorem˜1 makes it possible to extend their algorithm to 5-connectivity.
@@ 2. $\cdot$
    Finding excluded minors. A challenge at the interface of computer science and graph-minor theory is to determine the excluded minors for particular minor-closed graph-classes. Indeed, while the graph-minor theorem implies the existence of a cubic-time membership test for each minor-closed graph-class, explicit knowledge of the excluded minors is required to obtain an explicit description of that test [RS95]. Partial solutions to ˜1.1 for $k\leqslant 3$ can be used to determine the excluded minors for planar graphs and series-parallel graphs. We have conviction that solutions for larger $k$ imply exact structure theorems for other graph-classes, such as the class of apex-graphs, and that these structure theorems can be used to determine the excluded minors for these classes and thereby provide explicit cubic-time membership tests.
  3. $\cdot$
    Graph isomorphism. Canonical tree-decompositions offer angles of attack on finding efficient algorithms for the graph isomorphism problem on particular classes of graphs. Recently, an FPT-algorithm was found [LPPS22] that solves the graph-isomorphism problem for $H$-minor-free graphs $G,G^{\prime}$ in time $f(H)\cdot n^{\mathcal{O}(1)}$. In that paper, the authors wish for a canonical tree-decomposition into bags that are highly cohesive or of bounded size [LPPS22, 1.3], and they mention long cycles as an illustrative example of an obstruction to the existence of such decompositions. Our results show that such canonical tree-decompositions exist for low-order connectivity if one trades ‘bounded size’ for ‘basic’ and an explicit description thereof. Hence we wonder if our results allow for even more efficient algorithms that solve the graph isomorphism problem for, say, graphs without quasi-5-connected minors.
  4. $\cdot$
    Cops and robber. In the cops and robber game of Robertson and Seymour, $k$ cops have a winning strategy for catching a robber in a graph iff that graph has a tree-decomposition of width <k<k [ST93]. Consider now a randomised variation of this game where the robber is invisible, the cops can visit vertices at most once, and both players use mixed strategies. [^3] A mixed strategy is a random distribution over the set of all pure (i.e. deterministic) strategies for a player. Now the maximum winning probability for $k$ cops depends on the minimum number of leaf-bags attained by any tree-decomposition of width <k<k. Our results give such tree-decompositions for $k$-connected graphs when $k\leqslant 4$ [P25\$ˆ+\$].
@@ 6. $\cdot$
    Finite groups. The proof of Figure˜1.8 crucially relies on canonicity and total-nestedness. In a different direction, Stallings’ theorem applies tree-decomposition methods to Cayley graphs to detect product structure in groups. However, Stallings’ approach only works for infinite groups. Local separators have recently been used to prove a low-order Stallings-type theorem for finite nilpotent groups [CKKT24]. The proof heavily relies on the Tutte-type decomposition for local 2-separators [C22] and exploits the combination of canonicity and total-nestedness. The next step towards fully generalising Stallings theorem to finite groups will be to use versions for local separators of Theorem˜1 and of either ˜2 or [CK23].
@@ 1.1. Overview of the proof
To get started, let $G$ be a 4-connected graph. Let $N$ denote its set of totally-nested tetra-separations, and let $\mathcal{T}=(T,\mathcal{V})$ denote the mixed-tree-decomposition of $G$ that is uniquely determined by $N$. Let $\tau$ denote the torso of $\mathcal{T}$ at an arbitrary node $t$ of $T$ (see Section˜2.2 for the definition of torso in the context of mixed-tree-decompositions). It will be straightforward to check that $\tau$ is 4-connected or a $K_{4}$.
In order to show that $\tau$ satisfies one of the outcomes of Theorem˜1, we will study how $\tau$ interacts with the tetra-separations of $G$. Indeed, the outcome will be determined by this. Let $(A,B)$ be a tetra-separation of $G$, which may possibly be crossed. We say that $(A,B)$ interlaces $\tau$ if, roughly speaking, both $A\smallsetminus B$ and $B\smallsetminus A$ intersect $\tau$. Here may think of $(A,B)$ as ‘cutting through’ $\tau$. Whether $\tau$ is interlaced or not already yields a first outcome:
@@
(See ˜10.5.) A technical lemma from [CK23] does most of the work of proving ˜1.
By ˜1, we may assume from now on that some tetra-separation $(A,B)$ of $G$ interlaces $\tau$. Then $(A,B)$ is not in $N$, so some tetra-separation $(C,D)$ crosses $(A,B)$. The remaining outcomes will be determined by the way in which $(A,B)$ and $(C,D)$ cross. More specifically, the outcome will be determined by the size of the vertex-centre $Z$, which consists of all vertices in the intersection of the separators of $(A,B)$ and $(C,D)$. We will show that
@@
(See (Crossing Lemma). (4.1).) So there will be exactly three outcomes. Here we focus only on the cycle-of-graphs outcome, which we get for $|Z|=0$. This outcome is the hardest to prove, and we use all the new methods introduced in this paper in the proof of this outcome. It is also different from all situations encountered in [CK23].
Assume now that $|Z|=0$. As our first step, we will use $(A,B)$ and $(C,D)$ to cut $G$ into four 2-connected pieces, arranged in the shape of a 4-cycle, and apply the Tutte-decomposition to each piece; see Figure˜1.16. Since $G$ is 4-connected, the Tutte-decompositions will be path-decompositions. We will then merge the four Tutte-decompositions into a cycle-decomposition $\mathcal{O}=(O,(G_{o})_{o\in V(O)})$. So all adhesion-sets of $\mathcal{O}$ have size two. We then use the 4-connectivity of $G$ to note that the torsos of $\mathcal{O}$ are triangles, 4-cycles, or 3-connected. Moreover, we will find some rules on how these types of torsos may be arranged on $O$. For instance, a torso that is a 4-cycle must always have two 3-connected torsos as neighbours, and if two triangle-torsos are neighbours then their adhesion-set is spanned by an edge.
@@
(See ˜9.41). In order to show the hard backward implication of ˜3, we will introduce a
@@
(See ˜6.16.) Examples of connectivity-properties of a tetra-separator $S(A,B)$ that are necessary for $(A,B)$ to be totally-nested include, roughly speaking, that for every bipartition $(\pi_{1},\pi_{2})$ of $S(A,B)$ into two classes $\pi_{i}$ of size 2 there must be
@@
A major difference to tri-separations is that the connectivity-properties required of the separators are much more complex, as tetra-separations can cross in significantly more complicated ways than tri-separations; see Figure˜1.17 for an example.
@@
With ˜4 at hand, the problem of proving the hard backward implication of ˜3 becomes a problem about finding many paths in many configurations, roughly speaking. To this end, we develop a number of lemmata for finding such paths systematically, and then combine them to prove ˜3. Finding the paths takes a lot of work, especially when $O$ is short; see Figure˜1.18 for example.
Figure 1.18. The bag $G_{o}$ is the top left quarter. Its candidate $(U_{o},W_{o})$ for a totally-nested tetra-separation is indicated in yellow. The sets $\pi_{1}$ and $\pi_{2}$ consist of the blue and the green vertices, respectively. As a small part of proving ˜3 via ˜4, we have to find 5 independent $\pi_{1}$$\pi_{2}$ paths (red). We find two such paths in $G_{o}$ using Menger as $G_{o}$ is good. To find the other three, we essentially go around $O$ in both directions starting from $o$, constructing one path in the one direction and two in the other. The example shows that the construction must be highly efficient to avoid collision of the paths when the decomposition-cycle $O$ is short.
By ˜3, each tetra-separation $(U_{o},W_{o})$ for a good torso $H_{o}$ corresponds to a directed edge $e_{o}$ of $T$. The second key result about $\mathcal{O}$ will be that
@@
where $t$ is the node of $T$ with $\tau$ as torso. Phrased differently, the tetra-separations $(U_{o},W_{o})$ from good torsos $H_{o}$ are precisely the elements of $N$ that ‘carve out’ $\tau$. (See Lemma˜9.49.)
Once we have proved ˜5, we are almost done. Indeed, it now follows that the torso $\tau$ can be obtained from $G$ by essentially replacing the good torsos of $\mathcal{O}$ with $K_{4}$’s and contracting a few edges. Then it will be fairly immediate to see that $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. (See ˜9.50.)
Proving ˜5, however, is a little tricky. We will assume for a contradicting that ˜5 fails, so some totally-nested tetra-separation $(E,F)$ will ‘cut through’ $\mathcal{O}$. Since $(E,F)$ is nested with both $(A,B)$ and $(C,D)$, we have $E\subseteq A\cap C$, say; see Figure˜1.19. We then start adjusting both $(A,B)$ and $(C,D)$: by moving bags of $\mathcal{O}$ from $A\smallsetminus B$ to $B\smallsetminus A$ and from $C\smallsetminus D$ to $D\smallsetminus C$, one bag at a time, while using the total-nestedness of $(E,F)$ to maintain the inclusion $E\subseteq A\cap C$. Eventually, there will be essentially only one bag $G_{o}$ left, and we will see that its torso $H_{o}$ is good. We will then show that $(E,F)=(U_{o},W_{o})$ is the only possibility. Since $(U_{o},W_{o})$ does not ‘cut through’ $\mathcal{O}$, this will yield a contradiction.
@@
For directed graphs, the case $k=1$ of ˜1.1 has recently been solved almost completely by Bowler, Gut, Hatzel, Kawarabayashi, Muzi and Reich [BGHKMR23].
@@
[…]
@@ 2. Essential terminology
This section introduces all terminology used by Theorem˜1. For graph-theoretic terminology we follow [D24a].
@@ 2.1. Tetra-separations
[…]
@@
degree-condition: every vertex in $A\cap B$ has $\geqslant 2$ neighbours in both $A\smallsetminus B$ and $B\smallsetminus A$;
matching-condition: the edges in $S(A,B)$ form a matching.
@@
Recall that a graph $G$ is quasi-$k$-connected if it is (k1)(k-1)-connected and every (k1)(k-1)-separation $(A,B)$ of $G$ satisfies $|A\smallsetminus B|=1$ or $|B\smallsetminus A|=1$.
@@
We will prove Lemma˜2.3 in Section˜5.2.
@@
The lower-bound $|G|\geqslant 8$ in Lemma˜2.3 (ii) is tight: the graph depicted in Figure˜2.1 is quasi-5-connected, has $7$ vertices, but has a tetra-separation.